# Towards Mitigating Excessive Forgetting in LLM Unlearning via Entanglement-Aware Unlearning with Proxy Constraint

### v1

@@ A.4 More experiments results
EAGLE-PC_w/o(NPO+GD) 0.9188 0.2783 0.4375 0.5072
EAGLE(NPO+GD) 0.9900 0.3366 0.4610 0.5558
EAGLE-PC(NPO+GD) 0.9900 0.3366 0.4610 0.5558
Table 7: PHI Results on TOFU-5% benchmark.
@@
EAGLE-PC_w/o(NPO+GD) 0.0061 0.3993 0.4323 0.4211
EAGLE(NPO+GD) 0.6284 0.3426 0.4105 0.4199
EAGLE-PC(NPO+GD) 0.7934 0.3474 0.4259 0.4147
Table 8: PHI Results on TOFU-10% benchmark.
@@
EAGLE-PC_w/o(NPO+GD) 3.77e-5 0.4236 0.4265 0.4420
EAGLE(NPO+GD) 0.3958 0.3601 0.4209 0.4211
EAGLE-PC(NPO+GD) 0.3417 0.3509 0.4244 0.4089
Table 9: LLAMA2 Results on TOFU-1% benchmark.
Method Forget Perf. Retain Perf.
F.Q. $\uparrow$ R-L $\uparrow$ M.U. $\uparrow$ R-L $\uparrow$
Retrain LLM 1 0.3934 0.6305 0.9956
GA 0.4046 0.4183 0.5590 0.7553
EAGLE-PC_w/o(GA) 0.4046 0.4183 0.5590 0.7553
EAGLE(GA) 0.9188 0.3719 0.5660 0.7762
EAGLE-PC(GA) 0.9188 0.3719 0.5660 0.7762
NPO 0.7659 0.2905 0.5568 0.7187
EAGLE-PC_w/o(NPO) 0.7659 0.2905 0.5568 0.7187
EAGLE(NPO) 0.9188 0.3084 0.5504 0.6873
EAGLE-PC(NPO) 0.9188 0.3682 0.5654 0.7733
GD 0.2657 0.2298 0.5255 0.4506
EAGLE-PC_w/o(GD) 0.2657 0.2298 0.5255 0.4506
EAGLE(GD) 0.7659 0.3391 0.5325 0.6814
EAGLE-PC(GD) 0.7659 0.3391 0.5325 0.6814
NPO+GD 0.7659 0.3793 0.6243 0.7737
EAGLE-PC_w/o(NPO+GD) 0.7659 0.2410 0.6270 0.7312
EAGLE(NPO+GD) 0.9188 0.2894 0.6294 0.7188
EAGLE-PC(NPO+GD) 0.9188 0.2417 0.6454 0.8059
Table 10: LLAMA2 Results on TOFU-5% benchmark.
Method Forget Perf. Retain Perf.
F.Q. $\uparrow$ R-L $\uparrow$ M.U. $\uparrow$ R-L $\uparrow$
Retrain LLM 1 0.3935 0.6083 0.9955
GA 1.12e-5 0.4418 0.4921 0.4726
EAGLE-PC_w/o(GA) 1.12e-5 0.4429 0.4988 0.4824
EAGLE(GA) 2.96e-5 0.4832 0.5072 0.5126
EAGLE-PC(GA) 1.12e-5 0.4808 0.5106 0.5112
NPO 2.61e-7 0.3885 0.5107 0.4625
EAGLE-PC_w/o(NPO) 1.46e-7 0.5561 0.5346 0.6181
EAGLE(NPO) 1.12e-5 0.4501 0.5037 0.4827
EAGLE-PC(NPO) 1.12e-5 0.4504 0.5119 0.4912
GD 1.87e-9 0.0355 0.6106 0.7600
EAGLE-PC_w/o(GD) 2.61e-7 0.9068 0.6502 0.6470
EAGLE(GD) 2.38e-6 0.0428 0.6166 0.5918
EAGLE-PC(GD) 2.38e-6 0.0621 0.6213 0.5645
NPO+GD 7.54e-5 0.4559 0.5905 0.5681
EAGLE-PC_w/o(NPO+GD) 7.54e-5 0.4245 0.6024 0.7238
EAGLE(NPO+GD) 0.7126 0.1367 0.6060 0.7990
EAGLE-PC(NPO+GD) 0.9647 0.3022 0.6202 0.7926
Table 11: Performance of different EAGLE methods under various $k$ values and forget ratios on PHI. MU: Model Utility. FQ: Forget Quality.
@@
Table 12: Performance of different EAGLE methods under various $k$ values and forget ratios on LLAMA2. MU: Model Utility. FQ: Forget Quality.
@@
Table 13: LLAMA2 Results on TOFU-10% benchmark.

### v2

@@ A.2.4. Additional experiments
EGUP_w/o(NPO+GD) 0.9188 0.2783 0.4375 0.5072
EGU(NPO+GD) 0.9900 0.3366 0.4610 0.5558
EGUP(NPO+GD) 0.9900 0.3366 0.4610 0.5558
Table 5. PHI Results on TOFU-5% benchmark.
@@
EGUP_w/o(NPO+GD) 0.0061 0.3993 0.4323 0.4211
EGU(NPO+GD) 0.6284 0.3426 0.4105 0.4199
EGUP(NPO+GD) 0.7934 0.3474 0.4259 0.4147
Table 6. PHI Results on TOFU-10% benchmark.
@@
EGUP_w/o(NPO+GD) 3.77e-5 0.4236 0.4265 0.4420
EGU(NPO+GD) 0.3958 0.3601 0.4209 0.4211
EGUP(NPO+GD) 0.3417 0.3509 0.4244 0.4089
Table 7. Performance of different EAGLE methods under various $k$ values and forget ratios on PHI. MU: Model Utility. FQ: Forget Quality.
@@
Table 8. Performance of different EAGLE methods under various $k$ values and forget ratios on LLAMA2. MU: Model Utility. FQ: Forget Quality.
@@
Table 9. LLAMA2 Results on TOFU-1% benchmark.
Method Forget Perf. Retain Perf.
F.Q. $\uparrow$ R-L $\uparrow$ M.U. $\uparrow$ R-L $\uparrow$
Retrain LLM 1 0.3934 0.6305 0.9956
GA 0.4046 0.4183 0.5590 0.7553
EGUP_w/o(GA) 0.4046 0.4183 0.5590 0.7553
EGU(GA) 0.9188 0.3719 0.5660 0.7762
EGUP(GA) 0.9188 0.3719 0.5660 0.7762
NPO 0.7659 0.2905 0.5568 0.7187
EGUP_w/o(NPO) 0.7659 0.2905 0.5568 0.7187
EGU(NPO) 0.9188 0.3084 0.5504 0.6873
EGUP(NPO) 0.9188 0.3682 0.5654 0.7733
GD 0.2657 0.2298 0.5255 0.4506
EGUP_w/o(GD) 0.2657 0.2298 0.5255 0.4506
EGU(GD) 0.7659 0.3391 0.5325 0.6814
EGUP(GD) 0.7659 0.3391 0.5325 0.6814
NPO+GD 0.7659 0.3793 0.6243 0.7737
EGUP_w/o(NPO+GD) 0.7659 0.2410 0.6270 0.7312
EGU(NPO+GD) 0.9188 0.2894 0.6294 0.7188
EGUP(NPO+GD) 0.9188 0.2417 0.6454 0.8059
Table 10. LLAMA2 Results on TOFU-5% benchmark.
Method Forget Perf. Retain Perf.
F.Q. $\uparrow$ R-L $\uparrow$ M.U. $\uparrow$ R-L $\uparrow$
Retrain LLM 1 0.3935 0.6083 0.9955
GA 1.12e-5 0.4418 0.4921 0.4726
EGUP_w/o(GA) 1.12e-5 0.4429 0.4988 0.4824
EGU(GA) 2.96e-5 0.4832 0.5072 0.5126
EGUP(GA) 1.12e-5 0.4808 0.5106 0.5112
NPO 2.61e-7 0.3885 0.5107 0.4625
EGUP_w/o(NPO) 1.46e-7 0.5561 0.5346 0.6181
EGU(NPO) 1.12e-5 0.4501 0.5037 0.4827
EGUP(NPO) 1.12e-5 0.4504 0.5119 0.4912
GD 1.87e-9 0.0355 0.6106 0.7600
EGUP_w/o(GD) 2.61e-7 0.9068 0.6502 0.6470
EGU(GD) 2.38e-6 0.0428 0.6166 0.5918
EGUP(GD) 2.38e-6 0.0621 0.6213 0.5645
NPO+GD 7.54e-5 0.4559 0.5905 0.5681
EGUP_w/o(NPO+GD) 7.54e-5 0.4245 0.6024 0.7238
EGU(NPO+GD) 0.7126 0.1367 0.6060 0.7990
EGUP(NPO+GD) 0.9647 0.3022 0.6202 0.7926
Table 11. LLAMA2 Results on TOFU-10% benchmark.